A Businessman interested in testing to see a difference in the average deals processed by the 3 adopted systems

a. report the descriptive statistics for each sample

b. report and analyze the ANOVA test for the 3 systems

Date Birmingham Memphis Little Rock Jackson
January 9 18 20 18 14
January 17 23 31 22 30
January 18 19 25 22 21
January 31 29 36 28 35
February 1 27 31 28 24
February 6 26 31 31 25
February 14 31 24 19 25
February 17 31 31 28 28
February 20 33 35 35 34
February 29 20 42 42 21
Answer #1

the descriptive statistics for each sample

Factor N Mean \bar{X}_{j} StDev s_{j} 95% CI
Birmingham 10 25.7 5.44 (21.56, 29.84)
Memphis 10 30.6 6.38 (26.46, 34.74)
Little Rock 10 27.3 7.47 (23.16, 31.44)
Jackson 10 25.7 6.36 (21.56, 29.84)

In this example, the hypotheses are:

  • H0: μ1 = μ2 = μ3 = μ4
  • H1: The means are not all equal.

Null hypothesis There is no significant difference in the the average deals processed by the 3 adopted systems among the four cities.
Alternative hypothesis At least one average deals processed by the 3 adopted systems is different among the four cities.
Significance level α = 0.05

Equal variances were assumed for the analysis.

Factor Information

Factor Levels Values
City 4 Birmingham AL, Memphis TN, Little Rock AR, Jackson MS

The test statistic for testing H0: μ1 = μ2 = ... = μ4 is:

F=\frac{\frac{\sum_{j=1}^{4}n_{j}(\bar{X}_{j}-\bar{X})^{2}}{3}}{\frac{\sum_{i=1}^{4}\sum_{j=1}{4}(X_{ij}-\bar{X}_{j})^{2}}{36}}=1.28

where

  • 27 = individual observation,
  • X = sample mean of the jth treatment (or group),
  • equation_image45.gif = overall sample mean,
  • k = the number of treatments or independent comparison groups, and
  • N = total number of observations or total sample size.

In order to determine the critical value of F we need degrees of freedom, df1=k-1 and df2=N-k. In this example, df1=k-1=4-1=3 and df2=N-k=40-4=36.

The critical value is F0.05,3,36 = 2.8662

and the decision rule is as follows: Reject H0 if F > 2.8662.

Analysis of Variance

Source DF Adj SS Adj MS F-Value P-Value
Factor 3 160.1 53.36 1.28 0.295
Error 36 1498.7 41.63
Total 39 1658.8

Conclusion:

We fail to reject H0 because F-value=1.28 < F-critical value= 2.8662 and also P-value=0.295> level of significance=0.05.

We have statistically no significant evidence at α=0.05

Hence, There is no significant difference in the the average deals processed by the 3 adopted systems among the four cities.